# Towards Understanding Bias in Synthetic Data for Evaluation

### v1

@@ 1. Introduction
2. Synthetic Test Collection Dataset
3. Human vs. Synthetic Query Analysis
@@ 4.1. Bland Altman Bias Analysis
Figure 1b presents a Bland-Altman plot222https://en.wikipedia.org/wiki/Bland-Altman_plot, which is used to visually assess the agreement between the judgments made by the LLM and human evaluators. The plot reveals a bias of approximately , indicating that, on average, the LLM tends to assign slightly higher relevance scores compared to human judges. This small negative bias suggests a consistent but minor tendency for the LLM to be more lenient in its scoring (i.e., tends to overestimate). The 95% Limits of Agreement, which range from approximately to , capture the expected variability in the differences between the LLM and human scores. These limits suggest that in most cases, the difference in scores between the LLM and the human evaluators lies within this range. The relatively wide span of the limits indicates substantial variability, implying that while the LLM’s judgments are often close to those of human evaluators, there are instances where they significantly diverge.
@@ 5. System Ranking Analysis
In this section, we analyse the bias that test collection may exhibit towards systems that are based on a similar language model to the one that was used in the synthetic test collection construction process. To do this, similar to (Rahmani et al., 2024a), we categorised the systems submitted to TREC DL 2023 based on the approach they use (i.e., language models used in their ranking or retrieval pipeline) using the metadata file released as part of SynDL333https://rahmanidashti.github.io/SynDL/ resource (Rahmani et al., 2024b). This results in four different system categories: systems based on GPT, T5, GPT + T5 (i.e., a combination of GPT and T5), and others (i.e., traditional methods such as BM25, or any model that does not use either GPT or T5).
Figure 3 shows that across all system types, we observe that system performance tends to be higher on synthetic queries compared to real queries, indicating that synthetic test collections, regardless of whether they are constructed using GPT-4 or T5, may be easier and thus overestimate system effectiveness. For example, in Figures 3b and 3f, where synthetic labels are compared to real queries under human labels, nearly all system types show a consistent upward shift, suggesting an overoptimistic view of their true performance. This pattern is particularly noticeable in GPT-only runs, which often appear above the diagonal, highlighting their tendency to benefit from synthetic setups.
@@ 6.1.1. Modeling
where indexes observations and indexes different retrieval runs. The term is a random intercept capturing variation across runs, assumed normally distributed with variance . Residual errors are also normally distributed with variance . is the score for observation . is a binary indicator for whether the judge is GPT-4 (reference: human judge). are dummy variables for system types (reference: traditional). is query difficulty, is query length by words, is average document length, indicates whether the query was generated by GPT-4, and is the number of models used in the system pipeline.

### v2

@@ 1. Introduction
2. Synthetic Test Collection Dataset
3. Human vs. Synthetic Query Analysis
@@ 4.1. Bland Altman Bias Analysis
Figure 1b presents a Bland-Altman plot, which is used to visually assess the agreement between the judgments made by the LLM and human evaluators. The plot reveals a bias of approximately , indicating that, on average, the LLM tends to assign slightly higher relevance scores compared to human judges. This small negative bias suggests a consistent but minor tendency for the LLM to be more lenient in its scoring (i.e., tends to overestimate). The 95% Limits of Agreement, which range from approximately to , capture the expected variability in the differences between the LLM and human scores. These limits suggest that in most cases, the difference in scores between the LLM and the human evaluators lies within this range. The relatively wide span of the limits indicates substantial variability, implying that while the LLM’s judgments are often close to those of human evaluators, there are instances where they significantly diverge.
@@ 5. System Ranking Analysis
In this section, we analyse the bias that a test collection may exhibit towards systems using language models similar to those employed during its construction. Following (Rahmani et al., 2024a), we categorised TREC DL 2023 systems by their underlying models using the SynDL metadata (Rahmani et al., 2024b). This yields four categories: GPT-based, T5-based, GPT + T5 (combined), and others (e.g., BM25 or models not using GPT/T5). Figure 3 shows that across all system types, we observe that system performance tends to be higher on synthetic queries compared to real queries, indicating that synthetic test collections, regardless of whether they are constructed using GPT-4 or T5, may be easier and thus overestimate system effectiveness. For example, in Figures 3b and 3f, where synthetic labels are compared to real queries under human labels, nearly all system types show a consistent upward shift, suggesting an overoptimistic view of their true performance. This pattern is particularly noticeable in GPT-only runs, which often appear above the diagonal, highlighting their tendency to benefit from synthetic setups.
@@ 6.1.1. Modeling
where indexes observations (i.e., queries) and indexes different retrieval runs. The term is a random intercept capturing variation across runs, assumed normally distributed with variance . Residual errors are also normally distributed with variance . is the score for observation . is a binary indicator for whether the judge is GPT-4 (reference: human judge). are dummy variables for system types (reference: traditional). is query difficulty, is query length by words, is average document length, indicates whether the query was generated by GPT-4, and is the number of models used in the system pipeline.